{F} The furniture company inexpensive tables and chairs. The production process of each is similar in the painting department each table take 4 hours of carpentry and 2 hours in the painting department. Each chair requires 3 hours of carpentry and 1 hour painting department. During the current product period 240 hours of carpentry time are available and 100 hrs in the painting is available. Each table sold yields of profit of $7 and chair produced sold for $5 profit. Find the best combination of table and chairs to manufacture in order to reach the maximum number of profit?


1
Expert's answer
2022-02-01T07:46:23-0500

Let "x_1" = number of tables

"x_2" = number of chairs

"z" = profit

Maximize "z=7x_1+5x_2"

Subject to constraints:


"4x_1+3x_2\le240" ( Carpentry constraints)


"2x_1+x_2\le100" (Painting constraints)


"x_1,x_2" "\ge0" ( Non-negativity constraints)


Initial system of equations


"-7x_1-5x_2+z=0\\4x_1+3x_2+s_1=240\\2x_1+x_2+s_2=100"


Where "s_1" and "s_2" are slack variables


"\begin{matrix}
 x_1& x_2&s_1&s_2&z&& \\
 4&3&1&0&0&240\\2&1&0&1&0&100\\-7&-5&0&0&1&0
\end{matrix}"


Pivot column is "1^{st}"

Test ratio:

"\frac{240}{4}=60" "\frac{100}{2}=50"


Pivot row is "2^{nd}"


Using Gaussian Elimination

"R_2÷2\to R_2\>" "-4R_2+R_1\to R_1"

"7R_2+R_3\to R_3"


"\begin{matrix}
 x_1&x_2&s_1&s_2&z&&\\
 0&1&1&-2&0&40\\1&0.5&0&0.5&0&50\\0&-1.5&0&3.5&1&350
\end{matrix}"


Pivot column is "2^\n{nd}"

Test ratio:

"\frac{40}{1}=40" "\frac {50}{0.5}=100"


Pivot row is "1^{st}"



Using Gaussian elimination

"-½R_1+R_2\to R_2\\1.5R_1+R_3\to R_3"


"\begin{matrix}
 x_1&x_2&s_1&s_2&z&&\\0&1&1&-2&0&40\\1&0&-0.5&0.5&0&30\\0&0&1.5&0.5&1&410
 
\end{matrix}"


"\therefore x_1=" Number of tables "=40"


"x_2=" Number of chairs"=30"


"z=" Profit"= \$ 410"